OpenStudy (gabylovesyou):

Connie needs to burn at least 300 calories at the gym today. She can burn 20 calories per minute on the treadmill and 10 calories per minute on the elliptical trainer. Which inequality below can be used to find the possible combinations of time she could spend on each machine and still meet her goal of burning at least 300 calories?

OpenStudy (gabylovesyou):

20t + 10e >= 300 20t - 10e >= 300 20t + 10e <= 300 20t - 10e <= 300

OpenStudy (gabylovesyou):

@jim_thompson5910

OpenStudy (anonymous):

Well first we can eliminate some choices. We know she wants to burn at LEAST 300 so our answer will be greater or equal to 300. We also know that if she's going to be doing both exercises then she'd be COMBINING them to get her fitness goal, only one answer displays that.

OpenStudy (gabylovesyou):

A?

OpenStudy (anonymous):

Yep.